OpenStudy (anonymous):

Can someone please help me? Factor the algebraic expression below in terms of a single trigonometric function. csc^2 x - 1

OpenStudy (anonymous):

you know csc^2(theta) = 1 + cot^2(theta) so csc^2 x - 1 = 1 + cot^2 x - 1 = cot^2 x +1-1 = cot^2 x

OpenStudy (anonymous):

That's it? Wow. I'm an idiot.